Parker Plastic, Inc., manufactures plastic mats to use with rolling office chairs. Its standard cost information for last year follows:

Standard Quantity Standard Price (Rate) Standard Unit Cost
Direct materials (plastic) 9 sq ft. $ 1.30 per sq. ft. $ 11.70
Direct labor 0.3 hr. $ 13.50 per hr. 4.05
Variable manufacturing overhead (based on direct labor hours) 0.3 hr. $ 1.30 per hr. 0.39
Fixed manufacturing overhead $619,080 ÷ 938,000 units) 0.66


Parker Plastic had the following actual results for the past year:

Number of units produced and sold 1,000,000
Number of square feet of plastic used 12,800,000
Cost of plastic purchased and used $ 15,360,000
Number of labor hours worked 336,000
Direct labor cost $ 4,200,000
Variable overhead cost $ 1,700,000
Fixed overhead cost $ 393,000

Calculate Parker Plastic’s direct labor rate and efficiency variances. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable.)

F / UF
Direct Labor Rate Variance      
Direct Labor Efficiency Variance
Answer #1

Labor rate variance = (SR-AR) * AH

= (13.50 - 4,200,000/336,000) * 336,000

= 336,000 F

Labor efficiency variance = (SH-AH)* SR

= (1,000,000*0.3 - 336,000) * 13.50

= 486,000 U

Comment if you face any issues